OpenStudy (anonymous):

Explain how to graph 2x + y = 14 using the x- and y-intercepts method. Be sure to use complete sentences and use the x- and y-intercepts method only.

OpenStudy (anonymous):

@hunterhoff substitute x=0 first into that equation. Tlel me what you get for the y-value.

OpenStudy (anonymous):

Tell*

OpenStudy (anonymous):

\[2(0)+y=14\]

OpenStudy (anonymous):

Is the y value 2?

OpenStudy (anonymous):

what's \[2 \times 0=\]

OpenStudy (anonymous):

\[2(0)=2 \times 0\]

OpenStudy (anonymous):

So what's 2 times 0 equal?

OpenStudy (anonymous):

Do you know your multiplication time tables?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good. So then the equation would be this right? \[0+y=14\]

OpenStudy (anonymous):

My internet connection is just a little slow sorry about that

OpenStudy (anonymous):

@hunterhoff

OpenStudy (anonymous):

right

OpenStudy (anonymous):

what's \[y+0=?\]

OpenStudy (anonymous):

fill in that question mark.

OpenStudy (anonymous):

0 ?

OpenStudy (anonymous):

Ahh huh. Now I found your porblem with maths. You don't know how to do addition and subtraction with pronumerals.

OpenStudy (anonymous):

problem*

OpenStudy (anonymous):

\[2x+0= 2x\] \[4z+0=4z\] \[z+0=z\] What's \[0+y=?\]

OpenStudy (anonymous):

y

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

Y

OpenStudy (anonymous):

Yes. Finally.

OpenStudy (anonymous):

Good job.

OpenStudy (anonymous):

So then now you got rid of the zero in the equation. What's y equal to now? That means what's the y-value.

OpenStudy (anonymous):

I dont know?

OpenStudy (anonymous):

\[0+y=14\] You don't remember your equation you specifically wrote it on this website so people can help you with it? memory loss?

OpenStudy (anonymous):

14???

OpenStudy (anonymous):

Excellent. Finally you made some way into this question.

OpenStudy (anonymous):

So....at the start of the question, didn't I tell you that I made x equal to zero so then I can find that y=14?

OpenStudy (anonymous):

so then you can find y=14*

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So WHEN x=0.............y=14. Right?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

And when you write out your coordinates or points on a graph. It's in this form right? \[(x, y)\]

OpenStudy (anonymous):

so that would be (0,14) :)

OpenStudy (anonymous):

Great. Now you found what's called the "y-intercept"

OpenStudy (anonymous):

Yayyy! :)

OpenStudy (anonymous):

The y-intercept is located at (0, 14). Now. Try and do the EXACT same thing when finding the x-intercept. This time, what do you make equal to zero?

OpenStudy (anonymous):

2 ?

OpenStudy (anonymous):

No. What's in your mind mate? Peanut butter and jelly or something? Where's the brain?!?!? I made x equal to zero so then I could FIND the Y-INTERCEPT. NOW. We are trying to FIND the X-INTERCEPT. What would I make equal to zero in order to find the X-INTERCET. I told you that I made x=0 to Find the Y-INTERCEPT. Now when you're trying to find the X-INTERCEPT what would you make ?=0. Fill in that question mark.

OpenStudy (anonymous):

INTERCEPT*

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

14=0?

OpenStudy (anonymous):

Y

OpenStudy (anonymous):

FINALLY. You cracked it.

OpenStudy (anonymous):

Sorry about that...

OpenStudy (anonymous):

so you make y=0.

OpenStudy (anonymous):

Do that to the equation. Just replace y with zero.

OpenStudy (anonymous):

Show me the end result of that.

OpenStudy (anonymous):

y

OpenStudy (anonymous):

so your final result when you replace y with zero is just \[y\]

OpenStudy (anonymous):

Wow...

OpenStudy (anonymous):

Your equation is \[2x+y=14\] You seem to have a very short memory loss of about 1 minute.

OpenStudy (anonymous):

You can't seem to put something in your head for more than a minute.

OpenStudy (anonymous):

You're replacing y with zero. Do that simple step and let me see the result when you replace y with zero.

OpenStudy (anonymous):

Omg... Now Im really confused!!! Im sorry... :( Is it 2x+0=14

OpenStudy (anonymous):

Finally!!!

OpenStudy (anonymous):

So what's \[2x+0=?\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

I don't think I'm capable of teaching you. It's really frustrating everytime you mess up big time. My hope is almost gone for you.

OpenStudy (anonymous):

ok... Thanks for trying, I really appreciate it...

OpenStudy (anonymous):

Is it (2,0)

OpenStudy (anonymous):

If not then dont bother

OpenStudy (anonymous):

7!!!! Its 7!!!! :D (7,0) right?

OpenStudy (anonymous):

it just took me a second to think about it! :D

OpenStudy (anonymous):

Dude...That took you...let's see...15 minutes...

OpenStudy (anonymous):

If you can't work these out, then you have to take a step back and do addition and subtraction and multiplication/division of a pronumeral.

OpenStudy (anonymous):

Thank you so much for helping me! :)

OpenStudy (anonymous):

That's just weak when you're doing a calculation of simple algebra in 15minutes!

OpenStudy (anonymous):

You need to practice, practice, practice,, practice, practice....and practice, practice, practice.....x1000000.

OpenStudy (anonymous):

Once you do that, you will be getting 100% in all maths exams.

OpenStudy (anonymous):

ok... thanks

OpenStudy (anonymous):

Right now, to be honest, your level of maths and what you're competing with in your grade, I reckon you will be in the remedial class if you don't practice your maths. Because you're not even getting above 50% if you're finding it hard to solve these questions. That's my honesty. I'm not adding cherries on top to make you feel better. That's being realistic about your level of math compared to your grade.